VIBRATIONALLY EXCITED C_{6}H

\underline{CARL A. GOTTLIEB}, M. C. McCARTHY, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138; Division of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138..

Following the detection of the linear carbon chain radical C_{6}H in space, many rotational lines in the X^{2}Π electronic ground state have been observed in a low pressure discharge at millimeter wavelengths^{a} and in a supersonic molecular beam at centimeter wavelengths^{b}. In the course of a laboratory search for new reactive hydrocarbon molecules with 6 carbon atoms, several series of harmonically related lines with rotational constants near that of C_{6}H in the ground vibrational state were observed in the millimeter-wave band. On the basis of the close agreement in rotational constants and intensities, two of the series were assigned to ^{2}Σ and ^{2}Δ states of a low-lying excited bending vibrational level of C_{6}H. The standard Hamiltonian with five spectroscopic constants reproduces the observed rotational spectrum of the ^{2}Δ state, but several high-order distortion terms in the spin-rotation interaction are needed to reproduce the spectrum of the 2Σ state of C_{6}H and C_{6}D. From the measured intensities of the rotational lines it appears that the ^{2}Σ state lies very close to ground but the ^{2}Δ state lies much higher in energy. A brief summary of the laboratory spectrum and applications to the astronomical observations will be presented.

aH. Linnartz, T. Motylewski, O. Vaizert, J. P. Maier, A. J. Apponi, M. C. McCarthy, C. A. Gottlieb, and P. Thaddeus, J. Mol. Spectrosc. \underline{197}, 1 (1999).

bM. C. McCarthy and P. Thaddeus, J. Chem. Phys. \underline{122}, 174308 (2005).